# Black holes immersed in dark matter: energy condition and sound speed

### v1

@@ I Introduction
Dark matter, which may cluster at the center of galaxies and close to BHs [23, 24], could significantly impact the dynamics of compact binaries and the way in which GWs or electromagnetic waves propagate [18, 25, 26, 8, 27, 28]. This prompts the question of how the existence of surrounding matter impacts the generation and transmission of GWs, as well as the electromagnetic characteristics of BHs. To answer this question, the knowledge of motivated matter distributions and corresponding spacetime geometry will be extremely useful.
@@
Nevertheless, recent efforts have emerged to enhance the analysis by incorporating certain relativistic effects, highlighting their potential to substantially influence the conclusions concerning detectability and parameter estimation. Ref. [39] has introduced spacetime geometry generated by a non-spinning BH within a core of matter, offering a first step towards studying BH physics in realistic dense environments for a specific choice of matter distribution. In Ref. [40] generic matter profiles were considered. They explore the influence of various families of DM halos on the geometry of BHs and their associated geodesic structures. Recently it has also been extended to wormholes [41]. These profiles have been used to compute the impact of DM profiles on the GW fluxes from extreme mass ratio inspirals immersed in such environment [39, 39, 40, 42]
Nonetheless, several aspects remain to be explored. Firstly, previous works assumed the radial pressure to be negligible. In our present study, we take into account a non-zero radial pressure. Secondly, it becomes evident that the matter structures examined in prior works and in our current work, exhibit unphysical sound speeds and, in specific instances, diverging anomalies. We address and mitigate some of these issues, conducting an examination of their effects on the matter distribution and as a consequence, on the geometry.
Figure 1: In the left column we show the behavior of the dominant energy condition with respect to the radius, down to $r=2.001$, for both the Einasto and Hernquist profiles. The condition is violated near the BH. In the right column, we show the behavior of tangential sound speed with respect to the radius for both the Einasto and Hernquist profiles. Near the BH the sound speed becomes faster than light. In the subplot, the region where the sound speed becomes unity is demonstrated.
@@ II General structure for individual distribution
We begin the work by taking established density distribution from the literature. Firstly, we take the following density structure in a similar manner as Ref. [40],
@@
The profile described in Eq. (1) has a quite generic structure allowing sufficient freedom to model the density by varying the parameters. A family of different profiles can be generated from it. However, we will not go into such details in the current work. We will rather consider a different profile, namely, the Einasto model (Ein) [44, 45, 46, 47]. The profile structure has a different structure from that of in Eq. (1), given by:
@@
with $n=6$, $d_{n}=3n-1/3+.0079/n$ [46, 47], and $\rho_{e}$ representing the density at the radius $r_{e}$, which defines a volume containing half of the halo mass. To make the paper easy to understand we will rename $r_{e}$ and $\rho_{e}$ as $r_{e}\to a_{0}$ and $\rho_{e}\to\rho_{0}$. In the current work, we will set the value of $\rho_{0}$ for all the density profiles by setting the total halo mass up to a certain radius. For this purpose, we use the following,
@@
Figure 2: In the first row the density and in the second row the pressure solution of the TOV equation is demonstrated. In the last row, the behavior of the dominant energy condition across the radius is demonstrated. Near the BH the condition gets violated. For a given $(a_{0},M_{Scale})$ both the C and NC solution is demonstrated. All the main plots are demonstrated from $r=2.1$ to $r=10^{6}$ while in the subplot up to $r=2.001$ is plotted.
@@ III.1 Effect of matter for vanishing radial pressure
Using the Energy momentum conservation equation the the solution for tangential pressure can be found to be,
@@
In Fig. 1 we plot $c_{st}^{2}$ and $p_{t}/\rho$ for Einasto and Hernquist profile. The dominant energy condition imposes $\rho\geq p_{t}$. As can be seen from the $p_{t}/\rho$ plot the dominant energy condition is violated in the near region for all the configurations. It is also noteworthy that in all the cases the sound speed becomes larger than the speed of light around $r\leq 3.5-5$, depending on the profile. This also explains the violation of the energy condition which is connected to the sound speed. This implies that the model starts becoming unphysical in nature near the BH. This problem is not related to the vanishing radial pressure as it persists even in the presence of the non-vanishing radial pressure. In the next section, we will investigate this aspect and how some of the nature of the problem changes if radial pressure is taken into account.
@@ III.2 Effect of matter for isotropic pressure
Given the mass function obtained above, one integrates the first equation in Eq. (10) for the pressure. Since the equation is first order, it requires the specification of a single boundary condition in order to provide a unique solution. At a large radius $r\sim 10^{7}$ we expect the pressure to be very small. However, due to a lack of knowledge of the equation of state of the matter, particular consistent values are unknown. To get a consistent value we will use the boundary condition that the value of the radial pressure at $r=10^{7}$ is exactly equal to the value of tangential pressure described by Eq. (7). Therefore we impose,
@@
$\begin{split}-\frac{dp}{dr}=&\frac{(\rho+p)(m(r)+4\pi r^{3}p)}{r(r-2m(r))}\\ \\ \frac{d\rho(r)}{dr}=&\mathcal{R}(r)\\ \frac{dm(r)}{dr}=&4\pi r^{2}\rho.\end{split}$ (14)
@@
The reason behind such imposition is to keep the sound speed subluminal as well as satisfy the virial theorem, i.e. $c_{s}^{2}\leq 1/3$, which will make the solution satisfy the dominant energy condition. However, this condition will create a kink in the sound speed profile which must be addressed in the future. The boundary condition of pressure is taken to be the same as before, while the boundary condition of density is taken such that at a large distance density profile matches that of Hernquist or Einasto,
$\begin{split}p(r=10^{7})=&\,p_{t}(r=10^{7}),\\ \rho(r=10^{7})=&\,\rho_{Ein/Hern}(r=10^{7}).\end{split}$ (16)
@@
Figure 4: In the above figures the solution of the metric is demonstrated. In the first row metric component $g_{tt}$ is shown. All the metric plots are demonstrated from $r=2.001$ to $r=10^{6}$ to demonstrate the structure for individual profiles. However, in the right subplots from $r=6$ to $r=10^{6}$ is shown and in the left subplots from $r=2.001$ to $r=2.1$ is shown. In the second row, the deviation function $f(r)$ is shown. All the deviation function plots are demonstrated from $r=2.05$ to $r=10^{6}$ while in the subplot from $r=2.001$ to $r=2.05$ is shown. In all the cases the metric goes to zero rapidly near $r\sim 2$. The deviation function increases near $r\sim 2$. The increase is more for the C solution compared to the NC solution. Despite this increase, the $g_{tt}$ component rapidly takes small value near $r\sim 2$. In the final row $1-g_{tt}g_{rr}$ has been plotted. For a Schwarzschild BH even near the horizon, $g_{tt}g_{rr}$ stays equal to unity. In the current case, although this factor deviates from the BH value of 1, the deviation is not extremely large.
@@
Figure 5: In the first row of the above plots we demonstrate how the dominant energy condition varies across the radius for the density profiles which vanishes near the BH. The condition is clearly violated in the near zone. In the last row, we show the behavior of the magnitude of the tangential sound speed. We show the magnitude as $c_{st}^{2}$ becomes negative below the divergence point. All the profiles show diverging behavior in the near BH region. This feature is unique to this kind of density profile where the density starts to decrease after reaching the maxima.
@@
Figure 6: In the first row we plot $c_{st}^{2}$ and we zoom in the diverging region. For all the profiles, below a certain radius $c_{st}^{2}$ becomes negative. In the second row, we plot the derivative of the density function with respect to the radius. We find that in most of the range of radius, the derivative is negative and in the near zone it becomes positive. The point when it becomes zero corresponds to the maxima of the density function. As a consequence $c_{st}^{2}$ blows up at that point and becomes negative for a smaller value of radius. This indicates that the flattening and vanishing structure of density is directly connected to the unphysical behavior of the sound speed.
@@
The solutions found in this way are labeled as corrected solution (C) while the solutions from the previous approach are labeled as not corrected solution (NC), where not corrected refers to the not corrected energy condition and sound speed. In Fig. 2 we plot both the C and NC solutions. We also plot $p/\rho$. As it can be noticed, $p/\rho$ violates the dominant energy condition for NC solutions, whereas for C solutions the energy condition is satisfied. Interestingly the density and pressure for C solutions increase more with decreasing radius compared to the NC solutions. Therefore a matter profile that is consistent with causality seems to prefer an increased density near a BH rather than a density profile that is flattened. This aspect will be discussed more in later sections.
@@ IV.1 Metric solutions
Figure 7: In the first row of the above we demonstrate the pressure profile found by solving the TOV equation with the scaled density solution. Pressure in the C solution shows a steeper growth compared to the NC solution. In the second row, we demonstrate the density profile for the same. NC solutions show the flattened and decaying structure of density. The C solutions demonstrate an increase in density profile which is significantly different from the NC solution.
@@ IV.2 Coordinate singularity makes radial derivatives diverge
For a vacuum BH, $r=2$ is known to host a coordinate singularity where $g_{tt}$ goes to zero and $g_{rr}$ diverges. From the metric solutions demonstrated earlier similar behavior is demonstrated in the presence of the matter also. It may therefore imply that $dp/dr$ should diverge due to the denominator in Eq. (10) at $r=2m(r)$. This as a result can possibly have an impact on the sound speed calculations. However, we would argue that the singularity in Eq. (10) is just a coordinate singularity. Hence, it is not physical like a vacuum BH space-time. One should note that as we are outside a BH we should expect $r$ co-ordinate to be pathological at $r\to 2m(r)$. However, if we take the equivalent of the Tortoise coordinate, $\frac{dr^{*}}{dr}=e^{-\nu(r)}$, the equation simplifies to,
@@
V Impact of near horizon flatness of density profiles
In the previous sections, we considered Einasto and Hernquist profiles and studied the matter distribution, energy conditions, sound speed, and metric structure. We also demonstrated that resolving the sound speed introduces an increased overdensity region near the black hole.
With both the Newtonian and relativistic analyses it has been found that the density profiles with a BH at their core exhibit a vanishing density at the horizon and develop a cusp with a lengthscale determined by the BH’s mass [24, 23]. The specific form of the profile determines the behavior of the profiles, which could have implications for accurately modeling the GW signals emitted by coalescing binaries [33]. To assess the impact of the vanishing density near BH we introduce rescaling the density profile according to $\rho(r)\rightarrow\rho(r)(1-2/r)$, following the results of [10, 40] (check Appendix B for a different scaling).
@@ V.1 Matter distribution
Figure 9: In the above figures the solution of the metric is demonstrated for density profiles that vanish near the BH. In the left column metric component $g_{tt}$ is shown. All the metric plots are demonstrated from $r=2.001$ to $r=10^{6}$ to demonstrate the structure for individual profiles. However, in the subplot from $r=6$ to $r=10^{6}$ is shown. In the right column, the deviation function $f(r)$ is shown. All the deviation function plots are demonstrated from $r=2.05$ to $r=10^{6}$ while in the subplot from $r=2.001$ to $r=2.05$ is shown. In all the cases the metric goes to zero rapidly near $r\sim 2$. The deviation function increases near $r\sim 2$. The increase is more for the C solution compared to the NC solution. Despite this increase, the $g_{tt}$ component rapidly goes to vanishingly small value near $r\sim 2$.
@@
With the density and mass function at hand, it is straightforward to compute tangential pressure and corresponding tangential sound speed (Eq. (7) and (8)) when radial pressure is assumed to be vanishing. In Fig. 5 in the first row we demonstrate $p_{t}/\rho$ which violates the dominant energy condition like before. In the second row, we demonstrate the behavior of $|c_{st}^{2}|$. We find a divergent nature that is unique to these density structures. In the first row of Fig. 6 we show $c_{st}^{2}$ near the divergent region. With decreasing radius sound speed starts to grow and diverge. For even smaller values of radius square of the sound speed becomes negative.
@@
The origin can be understood by noticing the final row in Fig. 6. In the final row, we plot $d\rho/dr$. Although in most of the region it stays negative, eventually it starts to go towards the positive direction. In this region, as the derivative is negative the $c_{st}^{2}$ stays positive. Eventually, $d\rho/dr$ vanishes when the density reaches the maxima and afterward, the derivative becomes positive rendering $c_{st}^{2}$ negative. At the point where the maxima is reached (say $r_{m}$) $c_{st}^{2}$ blows up as the $d\rho/dr$ is in the denominator. However as ${r\to r_{m^{+}}}$ the derivative is a negative number and as ${r\to r_{m^{-}}}$ it is a positive number (where $r_{m^{-}}=r_{m}-\epsilon$ and $r_{m^{+}}=r_{m}+\epsilon$ with $\epsilon\to 0$) making $c_{st}^{2}$ take respectively the positive and negative values. Therefore the divergence arises solely from the flatness of the density function. The negative values of $c_{st}^{2}$ arise because after $r=r_{m}$ the density starts to decrease. Divergent behavior is present for all the profiles considered. Similarly, for all the profiles the density derivative crosses through the zero value. It points towards the possibility that under the single component ideal fluid description of DM outside a BH density can not get flattened otherwise sound speed will become unphysical. It may also imply that the energy-momentum structure assumed is not appropriate for the DM. This needs further investigation.
@@
In Fig. 7 we demonstrate the C solution of the TOV equations along with the NC solutions of density and pressure. It is very prominent here that the C solutions do not allow a flattening and eventually decreasing density. On the contrary, it makes the density grow near the BH. It is understandable as the flattening of the density profile is directly connected with the diverging sound speed. Therefore the imposition of $30<c_{s}^{2}<1/$ 3 does not allow either the flattening or decreasing density profile.

### v2

@@ I Introduction
Dark matter, which may cluster at the center of galaxies and close to BHs [23, 24], could significantly impact the dynamics of compact binaries and how GWs or electromagnetic waves propagate [18, 25, 26, 8, 27, 28]. This prompts the question of how the existence of surrounding matter impacts the generation and transmission of GWs, as well as the electromagnetic characteristics of BHs. To answer this question, the knowledge of motivated matter distributions and corresponding spacetime geometry will be extremely useful.
@@
Nevertheless, recent efforts have emerged to enhance the analysis by incorporating certain relativistic effects, highlighting their potential to substantially influence the conclusions concerning detectability and parameter estimation. Ref. [39] has introduced spacetime geometry generated by a non-spinning BH within a core of matter, offering a first step towards studying BH physics in realistic dense environments for a specific choice of matter distribution. In Ref. [40, 41, 42] generic matter profiles were considered. They explore the influence of various families of DM halos on the geometry of BHs and their associated geodesic structures. Recently it has also been extended to wormholes [43]. These profiles have been used to compute the impact of DM profiles on the GW fluxes from extreme mass ratio inspirals immersed in such environment [39, 39, 41, 44]
Nonetheless, several aspects remain to be explored. Firstly, previous works assumed the radial pressure to be negligible. In our present study, we take into account a non-zero radial pressure. Secondly, it becomes evident that the matter structures examined in prior works and our current work, exhibit unphysical sound speeds and, in specific instances, diverging anomalies. We address and mitigate some of these issues, examining their effects on the matter distribution and as a consequence, on the geometry.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 1: In this figure we consider $p_{r}=0$. In the left column, we show the behavior of the dominant energy condition with respect to the radius, down to $r=2.001$, for both the Einasto and Hernquist profiles. The condition is violated near the BH. In the right column, we show the behavior of tangential sound speed with respect to the radius for both the Einasto and Hernquist profiles. Near the BH the sound speed becomes faster than light. In the subplot, the region where the sound speed becomes unity is demonstrated.
@@ II General structure for individual distribution
We begin the work by taking established density distribution from the literature. Firstly, we take the following density structure as Ref. [41],
@@
The profile described in Eq. (1) has a quite generic structure allowing sufficient freedom to model the density by varying the parameters. However, we will not go into such details in the current work. We will rather consider a different profile, namely, the Einasto model (Ein) [46, 47, 48, 49]. The profile has a different structure from that of in Eq. (1), given by:
@@
with $n=6$, $d_{n}=3n-1/3+.0079/n$ [48, 49], and $\rho_{e}$ representing the density at the radius $r_{e}$, which defines a volume containing half of the halo mass. To make the paper easy to understand we will rename $r_{e}$ and $\rho_{e}$ as $r_{e}\to a_{0}$ and $\rho_{e}\to\rho_{0}$. For a given density we define $M_{\rm Halo}$ as,
[…]
A density profile therefore is completely determined if $a_{0}$ and $\rho_{0}$ are known. $M_{\rm Halo}$ is then known in terms of these two parameters. However, we will choose $M_{\rm Halo}$ to be the free parameter of a profile rather than $\rho_{0}$. Hence, in the current work, we will set the value of $\rho_{0}$ for all the density profiles by setting the total halo mass up to a certain radius. For this purpose, we use the following,
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 2: In this figure we consider $p_{r}=p_{t}=p$. In the first row the density and in the second row the pressure solution of the TOV equation is demonstrated. In the last row, the behavior of the dominant energy condition across the radius is demonstrated. For a given $(a_{0},M_{Scale})$ both the C and NC solution is demonstrated. Near the BH the condition gets violated for NC solutions. However, the C solutions do not violate it. All the main plots are demonstrated from $r=2.1$ to $r=10^{6}$ while in the subplot down to $r=2.001$ is plotted.
@@ III.1 Effect of matter for vanishing radial pressure
Using the Energy momentum conservation equation, the solution for tangential pressure can be found to be,
@@
In Fig. 1 we plot $c_{st}^{2}$ and $p_{t}/\rho$ for Einasto and Hernquist profile. The dominant energy condition imposes $\rho\geq p_{t}$. As can be seen from the $p_{t}/\rho$ plot the dominant energy condition is violated in the near region for all the configurations. It is also noteworthy that in all the cases the sound speed becomes larger than the speed of light around $r\leq 3.5-5$, depending on the profile. This also explains the violation of the energy condition which is connected to the sound speed. This implies that the model starts becoming unphysical near the BH. This problem is not related to the vanishing radial pressure as it persists even in the presence of the non-vanishing radial pressure. In the next section, we will investigate this aspect and how some of the nature of the problem changes if radial pressure is taken into account.
@@ III.2 Effect of matter for isotropic pressure
Given the mass function obtained above, one integrates the first equation in Eq. (11) for the pressure. Since the equation is first order, it requires the specification of a single boundary condition to provide a unique solution. At a large radius $r\sim 10^{7}$ we expect the pressure to be very small. However, due to a lack of knowledge of the equation of state of the matter, particular consistent values are unknown. To get a consistent value we will use the boundary condition that the value of the radial pressure at $r=10^{7}$ is exactly equal to the value of tangential pressure described by Eq. (8). Therefore we impose,
@@
$-\frac{dp}{dr}=$ $\frac{(\rho+p)(m(r)+4\pi r^{3}p)}{r(r-2m(r))}$ (15)
$\frac{d\rho(r)}{dr}=$ $\mathcal{R}(r)$ (16)
$\frac{dm(r)}{dr}=$ $4\pi r^{2}\rho.$ (17)
@@
The reason behind such imposition is to keep the sound speed subluminal as well as satisfy the virial theorem, i.e. $c_{s}^{2}\leq 1/3$, which will make the solution satisfy the dominant energy condition. However, this condition will create a kink in the sound speed profile which must be addressed in the future. Note, that this is the key innovation of the current work. This approach is completely different from previous approaches. As will be demonstrated later this innovation makes the sound speed physical.
The boundary condition of pressure is taken to be the same as before, while the boundary condition of density is taken such that at a large distance density profile matches that of Hernquist or Einasto,
$p(r=10^{7})=$ $\,p_{t}(r=10^{7}),$ (19)
$\rho(r=10^{7})=$ $\,\rho_{Ein/Hern}(r=10^{7}).$ (20)
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 4: In the above figures the solution of the metric is demonstrated. In the first row metric component $g_{tt}$ is shown. All the metric plots are demonstrated from $r=2.001$ to $r=10^{6}$ to demonstrate the structure for individual profiles. However, in the right subplots from $r=6$ to $r=10^{6}$ is shown and in the left subplots from $r=2.001$ to $r=2.1$ is shown. In the second row, the deviation function $f(r)$ is shown. All the deviation function plots are demonstrated from $r=2.05$ to $r=10^{6}$ while in the subplot, from $r=2.001$ to $r=2.05$ is shown. In all the cases the metric goes to zero rapidly near $r\sim 2$. The deviation function increases near $r\sim 2$. The increase is more for the C solution compared to the NC solution. Despite this increase, the $g_{tt}$ component takes a small value near $r\sim 2$. In the final row $1-g_{tt}g_{rr}$ has been plotted. For a Schwarzschild BH even near the horizon, $g_{tt}g_{rr}$ stays equal to unity. In the current case, although this factor deviates from the BH value of 1, the deviation is not extremely large.
Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 5: Here we show the results of scaling different density profiles by $(1-2/r)$. In the first row of the above plots, we demonstrate how the dominant energy condition varies across the radius for the density profiles which vanishes near the BH. The condition is clearly violated in the near zone. In the last row, we show the behavior of the magnitude of the tangential sound speed. We show the magnitude as $c_{st}^{2}$ becomes negative below the divergence point. All the profiles show diverging behavior in the near BH region. This feature is unique to this kind of density profile where the density has a maxima.
Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 6: In the first row we plot $c_{st}^{2}$ and we zoom in the diverging region. For all the profiles, below a certain radius $c_{st}^{2}$ becomes negative. In the second row, we plot the derivative of the density function with respect to the radius. We find that in most of the range of radius, the derivative is negative and in the near zone it becomes positive. The point when it becomes zero corresponds to the maxima of the density function. As a consequence $c_{st}^{2}$ blows up at that point and becomes negative for a smaller value of radius. This indicates that the maxima of the density is directly connected to the unphysical behavior of the sound speed.
@@
The solutions found in this way are labeled as corrected solution (C) while the solutions from the previous approach are labeled as not corrected solution (NC), where not corrected refers to the not corrected energy condition and sound speed. In Fig. 2 we plot both the C and NC solutions. We also plot $p/\rho$. As it can be noticed, $p/\rho$ violates the dominant energy condition for NC solutions, whereas for C solutions the energy condition is satisfied. Interestingly the density and pressure for C solutions increase more with decreasing radius compared to the NC solutions. Therefore a matter profile that is consistent with causality seems to prefer an increased density near a BH rather than a decrease in density after reaching maxima near the BH. This aspect will be discussed more in later sections.
@@ IV.1 Metric solutions
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 7: In the first row of the above we demonstrate the pressure profile found by solving the TOV equation with the scaled density solution. Pressure in the C solution shows a steeper growth compared to the NC solution. In the second row, we demonstrate the density profile for the same. NC solutions show a maxima and decaying structure of density. The C solutions demonstrate an increase in density profile which is significantly different from the NC solution.
@@ IV.2 Coordinate singularity makes radial derivatives diverge
For a vacuum BH, $r=2$ is known to host a coordinate singularity where $g_{tt}$ goes to zero and $g_{rr}$ diverges. From the metric solutions demonstrated earlier similar behavior is demonstrated in the presence of the matter also. It may therefore imply that $dp/dr$ should diverge due to the denominator in Eq. (11) at $r=2m(r)$. This as a result can have an impact on the sound speed calculations. However, we would argue that the singularity in Eq. (11) is just a coordinate singularity. Hence, similar to the vacuum BH space-time it is also not physical singularity. One should note that as we are outside a BH we should expect $r$ co-ordinate to be pathological at $r\to 2m(r)$. However, if we take the equivalent of the Tortoise coordinate, $\frac{dr^{*}}{dr}=e^{-\nu(r)}$, the equation simplifies to,
@@
V Impact of near horizon maxima of density profiles
In the previous sections, we considered Einasto and Hernquist profiles and studied the matter distribution, energy conditions, sound speed, and metric structure. We also demonstrated that the physical sound speed condition introduces an increased overdensity region near the black hole.
With both the Newtonian and relativistic analyses it has been found that the density profiles with a BH at their core exhibit a vanishing density at the horizon and develop a cusp with a lengthscale determined by the BH’s mass [24, 23]. The specific mathematical form of the profile determines the detailed nature of the profiles (i.e. the slope, maximum value, etc.), which could have implications for accurately modeling the GW signals emitted by coalescing binaries [33]. To assess the impact of the vanishing density near BH we introduce rescaling the density profile according to $\rho(r)\rightarrow\rho(r)(1-2/r)$, following the results of [10, 41] (check Appendix B for a different scaling).
@@ V.1 Matter distribution
Figure 9: In the above figures the solution of the metric is demonstrated for density profiles that vanish near the BH. In the left column metric component $g_{tt}$ is shown. All the metric plots are demonstrated from $r=2.001$ to $r=10^{6}$ to demonstrate the structure for individual profiles. However, in the subplot, from $r=6$ to $r=10^{6}$ is shown. In the right column, the deviation function $f(r)$ is shown. All the deviation function plots are demonstrated from $r=2.05$ to $r=10^{6}$, while in the subplot from $r=2.001$ to $r=2.05$ is shown. In all the cases $g_{tt}$ goes to zero rapidly near $r\sim 2$. The deviation function increases near $r\sim 2$. The increase is more for the C solution compared to the NC solution. Despite this increase, the $g_{tt}$ component rapidly goes to vanishingly small value near $r\sim 2$.
@@
With the density and mass function at hand, it is straightforward to compute tangential pressure and corresponding tangential sound speed (Eq. (8) and (9)) when radial pressure is assumed to be vanishing. In Fig. 5 in the first row we demonstrate $p_{t}/\rho$ which violates the dominant energy condition like before. In the second row, we demonstrate the behavior of $|c_{st}^{2}|$. We find a divergent nature that is unique to these density structures. In the first row of Fig. 6 we show $c_{st}^{2}$ near the divergent region. With decreasing radius sound speed starts to grow and diverge. For even smaller values of radius, the square of the sound speed becomes negative.
@@
The origin can be understood by noticing the final row in Fig. 6. In the final row, we plot $d\rho/dr$. Although in most of the region it stays negative, eventually it starts to go towards the positive direction. In this region, as the derivative is negative the $c_{st}^{2}$ stays positive. Eventually, $d\rho/dr$ vanishes when the density reaches the maxima and afterward, the derivative becomes positive rendering $c_{st}^{2}$ negative. At the point where the maxima are reached (say $r_{m}$) $c_{st}^{2}$ blows up as the $d\rho/dr$ is in the denominator. However as ${r\to r_{m^{+}}}$ the derivative is a negative number and as ${r\to r_{m^{-}}}$ it is a positive number (where $r_{m^{-}}=r_{m}-\epsilon$ and $r_{m^{+}}=r_{m}+\epsilon$ with $\epsilon\to 0$) making $c_{st}^{2}$ take respectively the positive and negative values. Therefore the divergence arises solely from the maxima of the density function. The negative values of $c_{st}^{2}$ arise because after $r=r_{m}$ the density starts to decrease. Divergent behavior is present for all the profiles considered. Similarly, for all the profiles the density derivative crosses through the zero value. It points towards the possibility that under the single component ideal fluid description of DM, outside a BH density can not reach a maxima, otherwise, sound speed will become unphysical. It may also imply that the assumed energy-momentum structure is inappropriate for the DM. This needs further investigation.
@@
In Fig. 7 we demonstrate the C solution of the TOV equations along with the NC solutions of density and pressure. It is very prominent here that the C solutions do not allow a decreasing density. On the contrary, it makes the density grow near the BH. It is understandable as the maxima of the density profile is directly connected with the diverging sound speed. Therefore the imposition of $30<c_{s}^{2}<1/$ 3 does not allow either the maxima or the decrease in density after the maxima.